Lionel left Reedsville at 9:00 A.M. one day and drove to Fitzwilliam at an average rate of 80 km/h. At 10:00 A.M. that same day, Nancy left Reedsville and followed the same route. If both Lionel and Nancy arrived in Fitzwilliam at 3:00 P.M., at what average rate had Nancy traveled? How far is Reedsville from Fitzwilliam?
Answer by mananth(16949)

Lionel speed = 80 kph
Nancy speed = x kph
Distance = speed * time
Lionel speed = 80 kph
time =6 hours (9.00 to 3.00 pm)
Distance between towns = 80*6=480km
Nancy speed =x
distance = 480
time = 5 hours
Nancy speed = 480/5
=96 kph